A 8 kg cubic block of unknown material weighs 6.3 kg when it is lowered completely into a tube containing mercury ( Density = 13.6x103 kg/ m3). Determine the side of the cube
Answer #1

Tuafo W = mga fb (6.3) X9.8= 889.8-Sm x V cube x 9.8 mg 1.7= 1.3.6x103 x Vcube. v=q3 = 1.25x10-4 m3 a=0.osma 5cm